. Let x be the binomial random variable with n = 10 and p = .9

a. Does the Empirical rule work on this distribution for data that is within one, two or three standard deviations of the mean? Explain.

b. Does Tchebysheff’s theorem work on this distribution for data that is within 1, 1.5, 2 or 4 standard deviations of the mean? Explain.
Answer #1

Solution :- Given n=101 P=0.9 aq=1-P=1-0.9=01 U= np=10X0.959 ronpa Vioxoxote 0-9486 a) Yes, Empherical sale tworky on it q+ 1